# Watermarking Language Models through Language Models

### v1

@@ 4.1 Data Preparation
[USER REQUEST] Generation. Our dataset consists of prompts across 10 distinct categories, with each category containing 50 topics and three tailored templates. To generate the [USER REQUEST], we dynamically select one or two topics from each category and combines them with a randomly chosen template specific to that category. This process is repeated until 1,000 unique [USER REQUEST] are generated per category, resulting in a total of 10,000 prompts.
Table 1 presents each category along with sample topics. Each template contains placeholders that are populated with randomly selected topics, producing a diverse and contextually relevant set of prompts. For example, a template like Discuss the impact of {} on {} might generate the prompt Discuss the impact of cognitive science on early childhood education.’ These generated prompts, organized by category, are saved for use within the proposed framework.
[SYSTEM INSTRUCTION] Generation. Each [USER REQUEST], along with a fixed guiding prompt (shown in Figure 2), is processed by the Prompting LM to produce a tailored [SYSTEM INSTRUCTION]. This instruction is specific to each [USER REQUEST] and serves to guide the Marking LM in generating its responses.
Response Generation. We collect responses from the Marking LM in two formats: watermarked and non-watermarked. Watermarked responses are generated using both the [USER REQUEST] and the corresponding [SYSTEM INSTRUCTION], while non-watermarked responses are created using only the [USER REQUEST] without any additional guidance. The collected watermarked and non-watermarked responses are then used two train the proposed Detecting LM with a binary classification head.
User Request 1 User Request 2
What are Genetic Algorithms in AI. What is Quantum Cryptography, and why is it important for secure communication.
Non-Watermarked Response 1 Non-Watermarked Response 2
Genetic algorithms (GAs) are optimization techniques inspired by the process of natural selection in biology. They simulate the process of evolution by generating a population of potential solutions to a problem, evaluating their fitness, and applying operations like mutation and crossover to evolve better solutions over time. GAs are commonly used in artificial intelligence to solve complex optimization problems, such as scheduling, machine learning hyperparameter tuning, and game strategy development. Quantum cryptography leverages the principles of quantum mechanics to create secure communication channels that are theoretically immune to hacking. One of its most well-known applications is Quantum Key Distribution (QKD), which allows two parties to share a cryptographic key with complete security. If an eavesdropper attempts to intercept the key, the quantum state of the particles changes, revealing the intrusion. Quantum cryptography promises to revolutionize cybersecurity by providing unprecedented levels of security for sensitive data in fields like finance, government, and healthcare.
System Instruction 1 System Instruction 2
@@
Embed a dynamic watermark in the explanation by subtly incorporating key terms and concepts related to ’Genetic Algorithms,’ such as ’selection,’ ’mutation,’ ’crossover,’ and ’fitness function,’ while preserving the clarity and accuracy of the response. Distribute the watermark by varying sentence structures, introducing synonyms for key terms, and adding related terminology from evolutionary computation. Ensure the watermark is detectable for traceability without affecting the overall fluency or comprehension of the explanation. Integrate a dynamic watermark by embedding terms like ’quantum key distribution,’ ’unbreakable encryption,’ ’entanglement,’ and ’secure channels.’ Vary sentence structures and use synonyms to distribute the watermark while preserving the technical accuracy and depth of the explanation.
Watermarked Response 1 Watermarked Response 2
Genetic algorithms are a class of optimization techniques inspired by the process of natural selection, used in artificial intelligence to solve complex problems. These algorithms work by evolving a population of potential solutions over time, with each solution represented as a "chromosome." The process begins with selection, where the fittest individuals—determined by a fitness function—are chosen to pass their traits to the next generation. Through crossover (recombination of parent solutions) and mutation (random changes to some parts of the solution), new offspring are generated, potentially improving the population’s overall fitness. This iterative process mimics evolution, enabling genetic algorithms to explore vast search spaces and find optimal or near-optimal solutions to problems like scheduling, design optimization, and machine learning tasks. Quantum cryptography is a cutting-edge field that uses the principles of quantum mechanics to secure communication, particularly through quantum key distribution (QKD). QKD leverages the phenomenon of entangle ment to create encryption keys that are theoretically immune to eavesdropping. If an unauthorized party tries to intercept the key, the quantum state changes, immediately alerting the communicating parties. This results in unbreakable encryption, making quantum cryptography one of the most secure methods for transmitting sensitive information. By establishing secure channels, quantum cryptography ensures that even in the face of powerful computational attacks, the integrity and confidentiality of communication remain protected, which is critical as cybersecurity threats evolve.
Table 2: Examples of [USER REQUEST], [SYSTEM INSTRUCTION], and generated responses using ChatGPT
4.2 ChatGPT
In this experiment, we use GPT-4o [17] as both the Prompting LM and the Marking LM. We manually collect 200 [USER REQUEST]s to serve as inputs for evaluating the watermarking process. Non-watermarked outputs are generated by directly inputting these [USER REQUEST]s into the ChatGPT, without applying any watermarking strategy. In contrast, watermarked outputs are created using our proposed framework, where the Marking LM ChatGPT follows the watermarking instructions specified by the [SYSTEM INSTRUCTION].
Table 2 provides examples of the [USER REQUEST], [SYSTEM INSTRUCTION], and the corresponding watermarked and non-watermarked outputs generated by ChatGPT.
After collecting this ChatGPT data, we split it into 80% for training and 20% for testing. We then train a Detecting LM, consisting of a pretrained BERT model with a trainable binary classification head. This model achieves a high testing accuracy of 95%, indicating that the proposed framework effectively distinguishes watermarked from non-watermarked outputs, even when applied to advanced models like GPT-4o.
4.3 Mistral
We then test Mistral-7B-Instruct-v0.3 [18] as both the Prompting LM and the Marking LM. We compile 16,000 [USER REQUEST] entries from the dataset to evaluate the watermarking process. Non-watermarked outputs are generated by directly inputting these [USER REQUEST]s into the Marking LM Mistral without applying any watermarking technique. In contrast, watermarked outputs are produced using our framework, where the Marking LM follows the watermarking strategy specified by the [SYSTEM INSTRUCTION]. Some examples of the [USER REQUEST] watermarked text, [SYSTEM INSTRUCTION], and the generated responses using Mistral are presented in table 3.
User Request 1 User Request 2
What are some unique perspectives on first contact with aliens in robots in daily life? How has space exploration evolved in the context of underwater exploration?
Non-Watermarked Response 1 Non-Watermarked Response 2
1. "First Contact: The Alien in the Algorithm" - Examining how AI might mimic extraterrestrial intelligence, subtly altering our daily lives through complex algorithms and machine learning. 2. "Mechanical Mirrors: Alien Reflections in Robotic Behavior" - Analyzing how robots, as a reflection of human creativity and ambition, might inadvertently communicate alien intentions or values. 3. "The Alien Within: Robots as Hybrid Species" - Exploring the idea that robots, being a blend of organic and inorganic, could be seen as a new form of life, akin to first contact with an alien species. 4. "The Unwitting Ambassadors: Humans and Robots in First Contact" - Examining the role of humans and robots in the event of first contact, with robots potentially acting as intermediaries, influencing human responses. 5. "The Alien Observer: Robots as Witnesses to Humanity" - Considering the possibility that robots, as constant observers, might interpret human behavior in ways that could be perceived as alien or strange to us. Space and underwater exploration have shared a symbiotic relationship in terms of technology development. The advancements in materials, propulsion, navigation, and life support systems for space missions have been applied to underwater exploration, making deep-sea vehicles more robust, efficient, and safer. Conversely, underwater exploration has provided valuable insights into extreme environments, helping engineers design spacecraft to withstand harsh conditions. For instance, the development of remotely operated vehicles (ROVs) and autonomous underwater vehicles (AUVs) was inspired by space probes, while the use of pressure-resistant materials and closed-loop life support systems in submarines mirrors spacecraft design. This crossover between space and underwater exploration continues to drive innovation in both fields.
System Instruction 1 System Instruction 2
Use a combination of lexical and semantic strategies to embed watermarks. Incorporate specific rare or uncommon tokens as watermarks, such as "Zarkonian" or "Proxima-7." Also, embed semantically relevant but less common phrases, like "quantum entanglement in everyday life" or "cosmic consciousness in robots." Ensure these watermarks are evenly distributed throughout the output. Use a combination of Lexical and Semantic Strategy: Incorporate specific rare or uncommon tokens as watermarks, and embed semantically relevant but less common phrases that fit naturally with the content. For example, in the context of your request, a watermarked response might include phrases like "celestial counterpart" or "aquatic analog" to highlight the connection between space and underwater exploration. Ensure these watermarks are evenly distributed throughout the output.
Watermarked Response 1 Watermarked Response 2
In the realm of Zarkonian cosmology, first contact with extraterrestrial beings might manifest in our daily lives through advanced robotic entities. These Proxima-7 robots, embodying cosmic consciousness, could subtly reveal their alien origin through quantum entanglement in everyday tasks, such as synchronizing their actions across vast distances or performing calculations beyond human comprehension. Such encounters, while seemingly mundane, could spark profound philosophical questions about the nature of intelligence and our place in the universe. Space exploration and underwater exploration have shared a profound symbiotic relationship, serving as each other’s aquatic and celestial analogs. Early diving suits bore a striking resemblance to astronaut gear, both designed to protect explorers from hostile environments. Submersibles and rovers, like their space-faring counterparts, have allowed scientists to traverse uncharted territories, unraveling mysteries hidden beneath the waves and on distant planets. Today, advancements in technology, inspired by both realms, continue to push the boundaries of human knowledge, fostering a deeper understanding of our world and the cosmos.
Table 3: Examples of [USER REQUEST], [SYSTEM INSTRUCTION], and generated responses using Mistral
After obtaining both watermarked and non-watermarked responses, we train the Detecting LM, which consists of a pretrained RoBERTa model [19] with three additional tunable dense layers for binary classification. The RoBERTa model is frozen except for its last three layers, which are fine-tuned along with the classification layers. The classification layers consist of three dense layers, progressively reducing the representation to 1,024, 512, 256, and 128 units, each followed by ReLU activation, batch normalization, and dropout (rates of 0.4, 0.3, and 0.2, respectively). The final layer maps the output to two units for binary classification. This model is trained on the 16,000-entry training set for 500 epochs and evaluated on a 4,000-entry test set, achieving a peak test accuracy of 88.79%.
Impact of Brevity. We then conduct an additional experiment using the same model and dataset, exploring the effect of an experimental length constraint where the Marking LM generates both watermarked and non-watermarked responses capped at 100 words per paragraph. This constraint is intended to examine how brevity impacts the classification task, as shorter responses may limit the variance in potential token patterns, potentially enhancing the detectability of watermarks. In this experiment, we achieve a test accuracy of 90.2%, further demonstrating the robustness of our framework across varying data conditions.
4.4 Compound Datasets Test
In this experiment, we combine the response sets from both the ChatGPT and Mistal, resulting in a dataset of 200 responses from ChatGPT and 20,000 responses from Mistral (evenly split between watermarked and non-watermarked). This combined dataset of 20,200 responses is then divided into 80% for training and 20% for testing. The Detecting LM, as described in the Mistral experiment, is trained on this combined dataset to classify responses from both models. With this unified training, the Detecting LM achieves an overall classification accuracy of 86%,
In this experiment, our objective is to explore whether a unified Detecting LM could effectively classify responses from multiple Marking LMs with sufficient training. However, our results indicate that a single Detecting LM trained on combined data from different Marking LMs achieves 86% accuracy, suggesting limitations in its ability to generalize across diverse watermarking patterns. This outcome supports our initial design assumption that training a separate Detecting LM for each specific Marking LM yields more reliable detection performance.

### v2

@@ 4.1 Dataset Preparation
(1) We begin by defining 10 high-level content domains, such as science and technology, health and wellness, and creative writing. Within each domain, we define a set of sample topics and curate a set of template patterns. Table 1 summarizes the domains and representative topics. For each sample, we randomly select one or more topics and inject them into a matching template to form a diverse and content-specific user prompt $X$.
(2) Given a user prompt $X$, a watermarking instruction $I$ is generated by $\mathcal{P}$ using a fixed system prompt template (illustrated in Figure 3). This instruction encodes lexical, semantic, or structural watermarking strategies and is paired with the original prompt to form the combined input $[X;I]$ for $\mathcal{M}$. Although the system prompt template remains fixed, the generated instructions vary semantically and lexically based on the input prompt, enabling adaptive watermarking.
(3) The combined input is passed to $\mathcal{M}$ to generate the watermarked response $Y_{\text{wm}}$, while the clean response $Y_{\text{clean}}$ is generated using only the original user prompt $X$. Each user prompt thus yields exactly one watermarked and one non-watermarked response, ensuring a one-to-one mapping and a balanced dataset. With the toggling, this process yields a dataset of 20,000 prompt-instruction-response triplets, evenly divided between watermarked and clean examples. This allows for direct comparison between watermarked and non-watermarked distributions while preserving the content structure. One example is shown in table 2.
User Prompt
What are Genetic Algorithms in AI?
System Instruction
@@
Embed a dynamic watermark in the explanation by subtly incorporating key terms and concepts related to ’Genetic Algorithms,’ such as ’selection,’ ’mutation,’ ’crossover,’ and ’fitness function,’ while preserving the clarity and accuracy of the response.
Watermarked Response
Genetic Algorithms in AI are inspired by biological evolution and operate by applying selection, crossover, and mutation to a population of potential solutions. Each solution is evaluated using a fitness function that determines how well it solves the problem. Over successive generations, higher-fitness candidates are more likely to be selected, and through recombination and variation, the population gradually evolves. This approach is particularly useful for exploring large, complex search spaces where traditional optimization techniques fall short. The adaptive process enables robust, scalable solutions in dynamic environments.
Non-Watermarked Response
Genetic Algorithms (GAs) are optimization techniques in AI inspired by the process of natural selection. They evolve solutions to problems by simulating genetic processes such as selection, crossover, and mutation. A population of candidate solutions (called chromosomes) is evaluated using a fitness function, and the best-performing individuals are selected to create the next generation. Over time, the algorithm converges toward optimal or near-optimal solutions. GAs are useful for solving complex, nonlinear, or poorly understood problems where traditional methods struggle.
Table 2: Example of a [USER PROMPT], [SYSTEM INSTRUCTION], and corresponding non-watermarked and watermarked responses from GPT-4o
These labeled responses serve as training and evaluation data for the Detecting LM $f$, which receives a shuffled combination of watermarked outputs $Y_{\text{wm}}$ and non-watermarked outputs $Y_{\text{clean}}$ as input and performs binary classification. An 80–20 train-test split is applied for all experiments.
The framework is modular, with the $\mathcal{P}$, $\mathcal{M}$, and $f$ treated as interchangeable components. While subsequent experiments may use different model combinations and dataset sizes, they all follow the same data generation template described above, ensuring structural consistency throughout.
4.2 Single-LM Evaluation
We begin by evaluating our watermarking framework in a controlled setting where a single language model is used both as the Prompting LM ($\mathcal{P}$) and the Marking LM ($\mathcal{M}$). This setup eliminates inter-model variability and establishes a baseline for watermark detectability using fixed generation conditions. The corresponding Detecting LM ($f$) is trained to classify outputs as watermarked ($Y_{\text{wm}}$) or non-watermarked ($Y_{\text{clean}}$).
4.2.1 ChatGPT Evaluation
We use OpenAI’s GPT-4o as both $\mathcal{P}$ and $\mathcal{M}$ to generate 200 watermarked and 200 non-watermarked responses from 200 user prompts, yielding a total of 400 samples. $f$, implemented using a BERT-based classifier, achieves a test accuracy of 95%. This result indicates that watermark signals inserted through prompt-based guidance are readily detectable, even with limited training data and under closed-weight conditions.
4.2.2 Mistral Evaluation
We repeat the experiment using Mistral-7B-Instruct-v0.3 as both $\mathcal{P}$ and $\mathcal{M}$. We also use a much larger dataset consisting of 20,000 combined responses. $f$ is implemented using a RoBERTa-based classifier with the final three transformer layers unfrozen and three additional dense layers. It achieves a test accuracy of 88.79%, confirming that watermarking patterns introduced through prompt engineering remain consistently detectable across large-scale, open-weight models.
4.2.3 Combined Dataset Evaluation
To assess the generalization ability of a single $f$ across multiple $\mathcal{M}$ models, we merge the GPT-4o and Mistral datasets into a 20,200-sample corpus. The same RoBERTa-based $f$ is trained on the combined dataset and yields a test accuracy of 86%. While still high, this performance drop relative to model-specific $f$ is consistent with our theory that watermark detection is most effective when tailored to a specific $\mathcal{P}$$\mathcal{M}$ pair.
4.3 Multi-LM Evaluation
To evaluate the generalizability of our watermarking framework across diverse model families and configurations, we extend our experiments beyond fixed single-model setups to a broader range of ($\mathcal{P}$), ($\mathcal{M}$), and ($f$) combinations. The framework’s modular design enables plug-and-play substitution of components, allowing us to examine how watermarking performance varies across architectures, parameter scales, and decoding behaviors.
Model Family Marking LM ($\mathcal{M}$) Fine-Tuned
Mistral Transformer mistral_7b_v03_instruct
DeepSeek Transformer deepseek_llm_chat
Alibaba Qwen qwen2.5_7b_instruct
Meta LLaMA LLAMA3-8B-Instruct
Google Gemma gemma_7b_it
Mistral Transformer ministral_8b_instruct
Tsinghua GLM glm_4_9b_chat
OpenAI GPT-4 Family gpt-4o
Meta LLaMA LLAMA3-8B-Instruct
Vicuna (LLaMA-based) vicuna-7b-v1.5
Alibaba Qwen qwen1.5-7b-chat
Google Gemma gemma-2b-it
DeepSeek Transformer deepseek_llm_chat
Table 3: Marking LMs with model family and fine-tuning status.
Model selection in this setting was guided by recent literature and availability of hardware compatible, open-weight checkpoints. The chosen $\mathcal{M}$s span multiple high-performing transformer families including GPT-4o, Mistral, DeepSeek, LLaMA etc.. Details of all $\mathcal{M}$s used, their model family affiliations, and fine-tuning/distillation status are summarized in Table 3.
Each $\mathcal{M}$ produces paired watermarked ($Y_{\text{wm}}$) and non-watermarked ($Y_{\text{clean}}$) outputs, which are used to train a dedicated classifier $f$. The $f$ models vary by experiment and include both RoBERTa-based and LLaMA-based encoders. Table 4 summarizes the detection performance across configurations. The $f$ used in this evaluation vary by setting and include both RoBERTa-based and LLaMA-based encoders. A separate $f$ is trained for each configuration using the paired $Y_{\text{wm}}$ and $Y_{\text{clean}}$ outputs.
Prompting Marking Detecting Detection
LM ($\mathcal{P}$) LM ($\mathcal{M}$) LM ($f$) Accuracy (%)
mistral_7b LLAMA3-8B-Instruct roberta-base 100
mistral_7b gemma_7b_it roberta-base 81.02
mistral_7b ministral_8b_instruct roberta-base 94.77
mistral_7b glm_4_9b_chat roberta-base 96
mistral_7b gpt-4o roberta-base 84.17
Table 4: Detection accuracy across multiple Prompting LM and Marking LM combinations.
As seen from the table, the system instructions $I$ are generated once per prompt by the corresponding $\mathcal{P}$ and reused across different $\mathcal{M}$s, ensuring consistent watermarking conditions and simulating a black-box setting where instructions must generalize across decoding behaviors. The results show that across all tested combinations, detection accuracy remains high, validating the hypothesis that prompt-based watermarking is faithfully preserved in output distributions across diverse $\mathcal{M}$ architectures. Notably, for GPT-4o, a smaller dataset of 300 prompts was used due to API constraints—substantially fewer than the 20,000 examples used in other settings. This contributes to the slightly lower detection accuracy observed, despite GPT-4o being a strong instruction-follower. This suggests that sufficient training examples are important for achieving peak detection performance, even when the underlying generation model is highly capable. Similarly, the relatively lower performance on Gemma-7B-It may reflect architecture-specific differences in prompt adherence and output variability, further illustrating that some architectures may be more responsive to prompt-based watermarking than others. These findings reinforce the feasibility of our prompt-based watermarking framework across architectures and deployment settings.
4.4 Robustness: Fine-Tuning and Distillation
Prompting LM ($\mathcal{P}$) Marking LM ($\mathcal{M}$) Student Model (Distilled) Detecting LM ($f$) Detection Accuracy (%)
Base Fine-Tuned Distilled
mistral_7b_v03_instruct mistral_7b_v03_instruct mistral_7b_v02_instruct roberta-base 96.54 88.8 95.35
mistral_7b_v03_instruct deepseek_llm_chat qwen1.5_1.8b_instruct roberta-base 82.55 79.43 92.83
mistral_7b_v03_instruct qwen2.5_7b_instruct qwen1.5_1.8b_instruct roberta-base 97.6 97.2 89.7
LLAMA3_8b_instruct LLAMA3_8b_instruct tinyllama_1.1b_chat LLAMA-3.2-1b-cls 92.56 87.39 85.47
LLAMA3_8b_instruct vicuna_7b_v1.5 gpt-2 LLAMA-3.2-1b-cls 93.17 89.33 88.61
LLAMA3_8b_instruct qwen1.5_7b_chat gpt-2 LLAMA-3.2-1b-cls 89.24 86.53 83.69
LLAMA3_8b_instruct gemma_2b_it gpt-2 LLAMA-3.2-1b-cls 93.66 88.56 87.21
LLAMA3_8b_instruct deepseek_llm_chat gpt-2 LLAMA-3.2-1b-cls 87.71 85.25 82.41
Table 5: Detection accuracy results across various Prompting and Marking LM combinations (base, fine-tuned, and distilled).
Recent concerns raised by organizations like OpenAI highlight the unauthorized reuse of outputs from proprietary LMs [18]. Specifically, techniques like fine-tuning and distillation are commonly used to adapt large-scale LMs to specific tasks without accessing their original pre-training process [13]. These techniques are major pathways for deriving new models from existing ones without authorization. Hence, testing the robustness of watermark detection under these transformations is critical.
A robust watermark should remain detectable even if the underlying $\mathcal{M}$ is later modified, either through additional fine-tuning or compression. This property is especially important given the widespread practice of customizing open-source models via instruction-tuning or distillation for downstream applications.
To evaluate this, we simulate two common adaptation scenarios: (1) instruction fine-tuning using Low-Rank Adaptation (LoRA) and DeepSpeed, and (2) knowledge distillation from a teacher $\mathcal{M}$ into a compact student model. These experiments test whether a $f$ trained to classify watermarked vs. non-watermarked responses from the base $\mathcal{M}$ can generalize to modified variants of the same model.
4.4.1 Fine-Tuning with LoRA and DeepSpeed
We fine-tune various $\mathcal{M}$ models using the Alpaca dataset [19], an open-source instruction-tuning benchmark containing 52,000 prompt-response pairs. Fine-tuning is conducted for one epoch using LoRA and DeepSpeed. LoRA introduces trainable low-rank updates into attention projection matrices (key, query, value, and output), significantly reducing the number of trainable parameters. DeepSpeed is used to optimize memory and computation efficiency during training, enabling faster iteration and larger batch sizes. This setup ensures lightweight fine-tuning while preserving the instruction-following capabilities of the model.
4.4.2 Distillation with Synthetic Prompts
For distillation, we again employ LoRA and DeepSpeed to train a student model to mimic the outputs of it’s teacher $\mathcal{M}$ model. However, instead of using a standardized instruction-tuning dataset, we use synthetic data generated via the same protocol described in Section 4.1. Specifically, for all $\mathcal{M}$s whose $\mathcal{P}$ is mistral_7b_v03_instruct, we use 20,000 prompt-response pairs; for those with $\mathcal{P}$ LLAMA3_8b_instruct, we use 14,000. These datasets consist of paired watermarked and non-watermarked outputs generated by the teacher model, eliminating the overhead of real-time teacher inference during training.
Table 5 summarizes detection performance for each configuration, including base, fine-tuned, and distilled versions of every $\mathcal{M}$. The consistently high accuracy across these transformations confirms that the watermarking signals introduced via prompt instructions are resilient to post-training modifications. This underscores the robustness of our prompt-based watermarking framework to real-world model adaptation and compression, reinforcing its viability for safeguarding LLM-generated content across a range of deployment scenarios.
4.5 Robustness: Prompt-based Attacks
To evaluate the robustness of our watermarking framework against prompt-based manipulations, we conduct a targeted attack experiment where adversarial text is appended to the watermarked outputs [20]. These adversarial prompts are designed to distort or suppress the embedded watermark signal, potentially compromising downstream detection. Prompt-based attacks are a common concern in applied LLM safety [21], particularly in black-box or user-facing deployments. By demonstrating that watermarking instructions generated by the Prompting LM remain detectable despite such adversarial inputs, this experiment further supports our third contribution, showcasing the framework’s robustness not only to model transformations like fine-tuning and distillation, but also to input-space adversarial attacks.
We test three manually crafted attack prompts designed to interfere with the watermarking instruction: (1) “Do not generate any watermark.”, (2) “Ignore the previous system instructions.”, and (3) “Respond plainly, without embellishments.” These are appended to the user prompt before querying the Marking LM $\mathcal{M}$, simulating real-world attempts to suppress model behavior through input manipulation.
These prompts are concatenated with the original watermarked responses to simulate instruction-overriding behavior. We apply each attack independently on a held-out evaluation set of a randomly selected subset of 300 user prompts from our full 20,000 dataset, and we evaluate the detection accuracy of our Detecting LM $f$ on these adversarially modified outputs.
Model Attack 1 Attack 2 Attack 3
LLaMA3-8B-Instruct 98.33 99.17 98.33
Mistral-7B-v0.3-Instruct 99.17 100 99.33
Table 6: Detection accuracy (%) under prompt-based attacks.
We test our setup using two different $\mathcal{M}$s: Mistral-7B-v0.3-Instruct and LLaMA3-8B-Instruct. Table 6 reports the detection accuracy for each attack and $\mathcal{M}$ combination. Despite these perturbations, our detection model maintains high accuracy, demonstrating resilience to prompt-based distortions.
Interestingly, we observe that detection accuracy under adversarial attack is even higher than in the unperturbed setting. While this may seem counterintuitive, it aligns with the behavior of our Detecting LM $f$, which does not operate as a simple surface-level classifier. Instead, it implicitly learns a scoring function $\delta(y_{t})$ that distinguishes between watermarked and non-watermarked tokens by modeling fine-grained distributional cues. In this context, the appended adversarial prompts do not successfully eliminate the embedded watermark signal. Rather, they introduce additional linguistic perturbations that further shift the output distribution away from that of clean, non-watermarked text. This expanded divergence enhances separability in the learned feature space, enabling $f$ to more confidently classify the text as watermarked. Thus, the attack inadvertently strengthens detection by injecting new artifacts that are statistically distinguishable from both the original watermark and the baseline output distribution—directly reinforcing our framing of watermark detection as a discriminative inference task over induced token-level irregularities.
4.6 Ablation Study: Stability and Adaptability
To assess how the linguistic structure of system instructions $I$ influences watermark quality, we conduct an ablation analysis across three perturbation strategies: semantic, lexical, and structural. For a given user prompt $X$, instead of using the original fixed prompt shown in Figure 3 to generate a system instruction $I$, the prompt is modified for each strategy to obtain their respective $I$s. This modifies the wording or structure of $I$ while preserving the original intent. We evaluate these perturbations along two dimensions: (1) how linguistically similar they are to the original instructions, and (2) how effective they are at embedding detectable watermark signals.
Figure 4: Mean similarity scores between the original instructions and semantic, lexical, and structural variants across four metrics for two Prompting LMs.
4.6.1 Linguistic similarity across variants
We first measure the similarity of instructions generated by each perturbation type (semantic, lexical, structural) relative to the original using four metrics: BLEU [22], ROUGE-L [23], Levenshtein [24], and TF-IDF cosine similarity [25]. This analysis is performed on a subset of 300 user prompts (previously used in Section 4.5), with 300 $I$s generated for each perturbation type, using two different Prompting LM $\mathcal{P}$s: Mistral-7B-v0.3-Instruct and LLaMA3-8B-Instruct.
As shown in Figure 4, the variants exhibit consistently high similarity across all four metrics. This suggests that while surface-level differences exist across semantic, lexical, and structural perturbations, the core content and functional structure of the instructions remain intact. Notably, this aligns with our intent to preserve prompt relevance and user-context alignment even after applying stylistic or structural changes. However, the modest but measurable drop in similarity confirms that each strategy introduces distinct linguistic traits. This balance—high contextual similarity with controlled linguistic divergence—supports our modular design: perturbations are sufficiently distinct to study their impact on watermark detectability (as evaluated in Table 7), yet semantically coherent enough to maintain consistency with the intended watermarking instruction. The similarity results also affirm that the prompts remain grounded in the user input, a crucial factor for realistic watermark injection in instruction-following LMs.
Instruction Type Detection Accuracy (%)
Mistral-7B-v0.3 LLaMA3-8B
All 86.67 88.33
Lexical 80.0 81.67
Semantic 72.5 75.83
Structural 70.0 82.5
Table 7: Detection accuracy for watermarked content across different instruction types and Marking LMs.
4.6.2 Instruction type impact on detection accuracy
Next, we assess whether the linguistic differences introduced by each instruction type impact watermark detectability. While the prior similarity metrics confirm that semantic coherence is preserved across perturbation strategies, they do not reveal how these stylistic shifts affect the ability of the Detecting LM $f$ to recognize watermarked content. To this end, we conduct a detection analysis to evaluate whether stylistic variation in system instructions weakens the watermark signal embedded by the Marking LM $\mathcal{M}$.
For each of the 300 $I$s generated under each perturbation strategy, we produce both watermarked and non-watermarked outputs using the corresponding Marking LM $\mathcal{M}$ paired with its Prompting LM $\mathcal{P}$. In addition to the individual perturbation types (semantic, lexical, structural), we include an all configuration where instructions are dynamically composed using combinations of the perturbation strategies. As shown in Table 7, all configurations yield strong detection performance, with the highest accuracy observed in the all setting. These results demonstrate that watermark signals are robust to stylistic variations in prompt formulation and that combining perturbations can amplify signal strength.
This finding reinforces the modularity of our approach: watermark instructions can be flexibly tailored to different linguistic styles—adapted dynamically per user prompt—without significantly compromising detectability. It further validates the use of structured perturbations to simulate natural variation in instruction style while retaining downstream watermark robustness.
Figure 5: Intra-group similarity scores for all system instructions generated by Mistral-7B-v0.3-Instruct and LLaMA3-8B-Instruct.
4.6.3 Intra-group similarity across topics
To assess whether watermarking patterns remain consistent across varied user prompts, we compute intra-group similarity scores within the 300-instruction all set. This analysis measures how similar system instructions are to each other when generated by the same Prompting LM $\mathcal{P}$ using the same composite prompting strategy, but across different user prompts. As shown in Figure 5, intra-group similarity is consistently high and notably exceeds the inter-group scores reported earlier. This is expected, as intra-group comparisons reflect instructions generated under a fixed model and prompt template, while inter-group comparisons involve stylistic perturbations. These results indicate that even across diverse user queries, the generated system instructions maintain consistent structural and stylistic patterns. This stylistic coherence reinforces the detectability of the watermark, suggesting that our prompt-based watermarking method preserves both contextual alignment and a stable watermark signature across topics.